# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 6.2 Composed Natural Image Retrieval
Traditional image retrieval systems often rely on single visual or textual information to retrieve the target image. However, this way is limited by the representation ability of a single modality, which makes it difficult to deal with complex image retrieval problems that involve multiple attributes or concepts. The main goal of Composed Natural Image Retrieval (CNIR) is to enhance the retrieval capability by composing the textual description with the image content, enabling the system to understand and respond to more abstract and complex user requirements. For instance, users can upload a photo of their favorite landscape with a description (e.g., “same location but in the Autumn”), and the CNIR system will parse the content of the image with the textual description, and understand the semantic associations to return a matched image. This technological advancement allows image retrieval to go beyond content similarity and capture deeper conceptual compositions.
CNIR is widely applied in various real-world scenarios, which has significant application value and business potential. In the field of travel, users can quickly find other images of the same scenic spot in a specific environment by uploading photos of the desired location and adding conditions such as season, time, and weather, providing inspiration and practical information for travel planning. Scientists can use the technology to conduct interdisciplinary research, such as ecologists combining geographic location and species image features to search for images in relevant literature. Furthermore, in the field of cultural heritage preservation, artists can use CNIR to compare images of artifacts from different periods of history and assist in restoration works. This precise and personalized search not only improves search efficiency and accuracy but also increases user participation and satisfaction, bringing new service models and development opportunities for the wide application of retrieval tasks.
@@ 6.2.1 Benchmark Datasets
CIRR [42] is the first released dataset for the natural domain. It consists of a total of 36554 triples derived from 21552 real-life images from the popular natural language reasoning dataset NLVR2 [85]. Each triplet consists of real-life images and human-generated modified sentences, arranged in an 80%-10%-10% split between the train/validate/test. This dataset encompasses rich object interactions, which addresses the issues of overly narrow domains and the high number of false negatives in the FashionIQ dataset [82]. Each query image in the validation and test sets has only one target image, where the test is evaluated on an online platform.
CIRCO [16] is the first dataset for CIR with multiple ground truths collected from the COCO 2017 unlabeled set [27], which is specifically designed for ZS-CIR evaluation. It consists of a total of 1020 queries, 220 for the validation set, and 800 for the test set, with an average of 4.53 ground truths per query. In addition, it uses all the 120K images of COCO as the index set, thus providing significantly more distractors than the 2K images of the CIRR test set. Compared to CIRR [42], CIRCO employs a two-phase annotation strategy to ensure higher quality, reduced false negatives, and the availability of multiple ground truths. Thus, it is also extremely challenging. The test is similarly evaluated on an online platform.
@@
SynthTriplets18M [9] is a vast set of high-quality 18M triplet datasets synthesized using large-scale generative models such as OPT [86] and Stable Diffusion [87, 88]. Specifically, they use two strategies: the keyword-based generation (11.4M), and the LLM-based generation (7.4M). SynthTriplets18M is over 500 times larger than existing datasets and covers a diverse and extensive range of conditioning cases.
Birds-to-Words [89] consists of images of birds from iNaturalist combined with human-annotated paragraphs to describe the difference between these pairs of images. This dataset is characterized by “long” natural language queries with every one of 3,347 image pairs having on average 4.8 paragraphs, each describing the differences between the pair of birds in an average of 31.38 words. Birds-to-Words provides richer text descriptions in each example than any of the other datasets current study, although the number of examples is small.
MIT-States [90] has 60k images, and each comes with an object/noun label and a state/adjective label (such as “red tomato” or “new camera”). For nouns, it selected words that refer to physical objects, materials, and scenes. For adjectives, it selected words that refer to specific physical transformations. Then, for each adjective, if a clear antonym exists, it is paired with another antonym adjective in the list. Finally, there are 249 nouns and 115 adjectives, on average each noun is only modified by 9 adjectives it affords.
@@ 6.2.2 Results and Analysis
CIRPLANT21’ICCV ResNet152 LSTM $15.18$ $43.36$ $60.48$ $33.81$ $56.99$ $75.40$
CIRPLANT W/ OSCAR21’ICCV ResNet152 LSTM $19.55$ $52.55$ $68.39$ $39.20$ $63.03$ $79.49$
Combiner22’CVPR-D CLIP(RN50) CLIP $33.59$ $65.35$ $77.35$ $62.39$ $81.81$ $92.02$
CLIP4Cir122’CVPR-W CLIP(RN50) CLIP $38.53$ $69.98$ $81.86$ $68.19$ $85.64$ $94.17$
NEUCORE23’NIPS-W ResNet BiGRU $18.46$ $49.40$ $63.57$ $44.27$ $67.06$ $78.92$
ComqueryFormer23’TMM Swin-T BERT $25.76$ $61.76$ $75.90$ $51.86$ $76.26$ $89.25$
ARTEMIS22’ICLR ResNet50 BiGRU $16.96$ $46.10$ $61.31$ $39.99$ $62.20$ $75.67$
RankUn23’Arxiv CLIP(RN50) CLIP $32.24$ $66.63$ $79.23$ $61.25$ $81.33$ $92.02$
CLIP4Cir223’TOMM CLIP(RN50) CLIP $42.05$ $76.13$ $86.51$ $70.15$ $87.18$ $94.40$
SSN23’Arxiv CLIP(ViT-B/32) CLIP $43.91$ $77.25$ $86.48$ $71.76$ $88.63$ $95.54$
TG-CIR23’ACMMM CLIP(ViT-B/16) CLIP $45.25$ $78.29$ $87.16$ $72.84$ $89.25$ $94.13$
MCEMiner24’TIP ResNet152 LSTM $17.48$ $46.13$ $62.17$
CaLa-CLIP4Cir24’SIGIR $35.37$ $68.89$ $80.04$ $66.68$ $84.65$ $93.42$
CaLa-ARTEMIS24’SIGIR $47.37$ $79.33$ $88.17$ $76.02$ $90.29$ $96.19$
CaLa-BLIP2cir24’SIGIR $49.11$ $81.21$ $89.59$ $76.27$ $91.04$ $96.46$
BLIP4CIR24’WACV BLIP BLIP $40.17$ $71.81$ $83.18$ $72.34$ $88.70$ $95.23$
BLIP4CIR+Bi24’WACV BLIP BLIP $40.15$ $73.08$ $83.88$ $72.10$ $88.27$ $95.93$
SPIRIT24’TOMM CLIP(RN50) CLIP $40.23$ $75.10$ $84.16$ $73.74$ $89.60$ $95.93$
DQU-CIR24’SIGIR CLIP(ViT-H/14) $46.22$ $78.17$ $87.64$ $70.92$ $87.69$ $94.68$
BLIP4CIR224’TMLR BLIP(ViT-B) BLIP $44.70$ $76.59$ $86.43$ $75.02$ $89.92$ $95.64$
CompoDiff(ST18M & FT)23’TMLR CLIP(ViT-L/14) CLIP $21.30$ $55.01$ $72.62$ $58.82$ $77.60$ $88.37$
CompoDiff(ST18M & FT)23’TMLR CLIP(ViT-G/14) CLIP $32.39$ $57.61$ $77.25$ $67.88$ $85.29$ $94.07$
SDQUR24’TCSVT BLIP2 $53.13$ $83.16$ $90.60$ $79.47$ $91.74$ $96.63$
NSFSE24’TMM ResNet152 BiGRU $20.70$ $52.50$ $67.96$ $44.20$ $65.53$ $78.50$
MANME24’TCSVT ResNet50 BiGRU $18.27$ $48.02$ $63.23$ $42.43$ $64.89$ $77.93$
SPRC24’ICLR BLIP2(ViT-L) $51.96$ $82.12$ $89.74$ $80.65$ $92.31$ $96.60$
@@
PALAVAR22’ECCV CLIP(ViT-B/32) $16.62$ $43.49$ $58.51$
Pic2Word23’CVPR CLIP(ViT-L/14) $23.90$ $51.70$ $65.30$
SEARLE23’ICCV CLIP(ViT-B/16) $24.00$ $53.42$ $66.82$ $54.89$ $76.60$ $88.19$
SEARLE-XL-OTI23’ICCV CLIP(ViT-L/14) $24.87$ $52.31$ $66.29$ $53.80$ $74.31$ $86.94$
SEARLE-XL23’ICCV CLIP(ViT-L/14) $24.00$ $53.42$ $66.82$ $54.89$ $76.60$ $88.19$
MTCIR23’Arxiv CLIP(ViT-L/14) $25.52$ $54.58$ $67.59$ $55.64$ $77.54$ $89.47$
VQA4CIR(+CLIP4CIR)23’Arxiv $51.40$ $81.71$ $86.65$ $73.16$ $88.84$ $95.59$
VQA4CIR(+SPRC)23’Arxiv $54.00$ $84.23$ $91.85$ $82.07$ $93.45$ $97.08$
KEDs24’CVPR CLIP(ViT-L/14) $26.40$ $54.80$ $67.20$
PM24’Arxiv CLIP(ViT-L/14) $26.10$ $55.20$ $67.50$ $56.00$ $76.60$ $88.00$
FTI4CIR24’SIGIR CLIP(ViT-L/14) $25.90$ $55.61$ $67.66$ $55.21$ $75.88$ $87.98$
@@
MCL(OPT-6.7B)21’ICML $24.15$ $55.98$ $90.92$ $59.52$
MCL(Llama2-7B)21’ICML $26.22$ $56.84$ $91.35$ $61.45$
CompoDiff(ST18M)23’TMLR CLIP(ViT-L/14) $18.24$ $53.14$ $70.82$ $57.42$ $77.10$ $87.90$
CompoDiff(ST18M)23’TMLR CLIP(ViT-G/14) $26.71$ $55.14$ $74.52$ $64.54$ $82.39$ $91.81$
TransAgg23’BMVC BLIP $38.10$ $68.42$ $93.51$
CASE24’AAAI BLIP(ViT)+BERT $48.00$ $79.11$ $87.25$ $75.88$ $90.58$ $96.00$
VISTA24’ACL $76.10$ $75.70$
HyCIR(CC3M+synthetic)24’Arxiv CLIP $25.08$ $53.49$ $67.03$ $53.83$ $75.06$ $87.18$
HyCIR(CC3M+synthetic)24’Arxiv BLIP $38.28$ $69.03$ $79.71$ $66.79$ $84.79$ $93.06$
VDG(Human+COCO’se)24’CVPR BLIP $49.37$ $78.12$ $85.52$ $76.68$ $90.46$ $96.05$
VDG(Human+NLVR2’se)24’CVPR BLIP $50.96$ $80.15$ $86.86$ $77.45$ $90.65$ $96.10$
PTG(SPRC+PTG)24’Arxiv $36.40$ $66.10$
TABLE VI: Comparison of Different Methods on CIRCO Dataset
Method Backbone CIRCO
Visual Textual mAP@1 mAP@10 mAP@25 mAP@50
@@
PALAVAR22’ECCV CLIP(ViT-B/32) $4.61$ $5.32$ $6.33$ $6.80$
MTCIR23’Arxiv CLIP(ViT-L/14) $10.36$ $11.63$ $12.95$ $13.67$
Pic2Word23’CVPR CLIP(ViT-L/14) $8.72$ $9.51$ $10.64$ $11.29$
SEARLE-XL-OTI23’ICCV CLIP(ViT-L/14) $10.18$ $11.03$ $12.72$ $13.67$
SEARLE-XL23’ICCV CLIP(ViT-L/14) $11.68$ $12.73$ $14.33$ $15.12$
FTI4CIR24’SIGIR CLIP(ViT-L/14) $15.05$ $16.32$ $18.06$ $19.05$
CIReVL24’ICLR CLIP(ViT-B/32) $14.94$ $15.42$ $17.00$ $17.82$
CIReVL24’ICLR CLIP(ViT-L/14) $18.57$ $19.01$ $20.89$ $21.80$
CIReVL24’ICLR CLIP(ViT-G/14) $26.77$ $27.59$ $29.96$ $31.03$
LinCIR24’CVPR CLIP(ViT-L/14) $12.59$ $13.58$ $15.00$ $15.85$
LinCIR24’CVPR CLIP(ViT-H/14) $17.60$ $18.52$ $20.46$ $21.39$
LinCIR24’CVPR CLIP(ViT-G/14) $19.71$ $21.01$ $23.13$ $24.18$
iSEARLE-XL-OTI24’Arxiv CLIP(ViT-L/14) $11.31$ $12.67$ $14.46$ $15.34$
iSEARLE-XL24’Arxiv CLIP(ViT-L/14) $12.50$ $13.61$ $15.36$ $16.25$
RTD(SEARLE)24’Arxiv CLIP(ViT-B/32) $11.26$ $12.11$ $13.63$ $14.37$

### v2

@@ 6.2 Composed Natural Image Retrieval
Traditional image retrieval systems typically rely on either visual or textual inputs alone. However, single-modality approaches are limited in their ability to represent complex queries involving multiple attributes or concepts. Composed Natural Image Retrieval (CNIR) addresses this limitation by integrating both textual descriptions and image content, enabling systems to better understand abstract and nuanced user requirements. For example, users can upload an image of a favorite landscape along with a description such as “same location but in Autumn,” and the CNIR system will analyze both modalities to retrieve semantically relevant images. By enabling more accurate and personalized search experiences, CNIR significantly improves efficiency and user satisfaction. It also opens new opportunities for innovative services and the broader development of advanced image retrieval technologies.
@@ 6.2.1 Benchmark Datasets
CSS [20] is a dataset close to the earliest, which consists of 32K queries (16K for training and 16K for test), such as ’make yellow sphere small’ that serve as modification text for the images synthesized in a 3-by-3 grid scene. Although it is a relatively simple dataset, CSS has the benefit of facilitating carefully controlled experiments.
CIRR [21] is the first released dataset for the natural domain. It consists of a total of 36554 triples derived from 21552 real-life images from the popular natural language reasoning dataset NLVR2 [22]. Each triplet consists of real-life images and human-generated modified sentences, arranged in an 80%, 10%, 10% split between the train/validate/test. This dataset encompasses rich object interactions, which addresses the issues of overly narrow domains and the high number of false negatives in the FashionIQ [26]. Each query image in the validation and test sets has only one target image, and the test is evaluated on an online platform.
CIRCO [23] is the first dataset for CIR with multiple ground truths collected from the COCO 2017 unlabeled set [242]. It consists of a total of 1020 queries, 220 for the validation set, and 800 for the test set, with an average of 4.53 ground truths per query. In addition, it uses all the 120K images of COCO as the index set, thus providing significantly more distractors than the 2K images of the CIRR test set. Compared to CIRR [21], CIRCO employs a two-phase annotation strategy to ensure higher quality, reduced false negatives, and the availability of multiple ground truths. Thus, it is also extremely challenging. The test is similarly evaluated on an online platform.
MIT-States [25] has 60k images, and each comes with an object/noun label and a state/adjective label (such as “red tomato” or “new camera”). For nouns, it selected words that refer to physical objects, materials, and scenes. For adjectives, it selected words that refer to specific physical transformations. Then, for each adjective, if a clear antonym exists, it is paired with another antonym adjective in the list. Finally, there are 249 nouns and 115 adjectives; on average, each noun is only modified by 9 adjectives it affords.
Birds-to-Words [24] consists of images of birds from iNaturalist combined with human-annotated paragraphs to describe the difference between these pairs of images. This dataset is characterized by “long” natural language queries, with every one of 3,347 image pairs having on average 4.8 paragraphs, each describing the differences between the pair of birds in an average of 31.38 words. Birds-to-Words provides richer text descriptions in each example than any of the other datasets in the current study, although the number of examples is small.
The following are representative datasets constructed through data generation.
@@
SynthTriplets18M [98] is a vast set of high-quality 18M triplet datasets synthesized using large-scale generative models such as OPT [243] and Stable Diffusion [244, 245]. Specifically, they use two strategies: the keyword-based generation (11.4M), and the LLM-based generation (7.4M). SynthTriplets18M is over 500 times larger than existing datasets and covers a diverse range of conditioning cases.
Good4cir [147] generates two datasets via a three-stage synthetic pipeline: CIRR-R rewrites CIRR captions to produce 199,350 training triplets (28,225 pairs) and 22,620 validation triplets (4,184 pairs). Captions avoid simplistic edits (e.g., ”add a red ball”) and instead offer fine-grained object modifications. Hotel-CIR: Mines hotel image pairs using perceptual hashing and CLIP embeddings, yielding 415,447 training triplets (65,364 pairs), 13,298 validation triplets (2,092 pairs), and 13,178 test triplets (2,069 pairs + 5 distractors/pair). Generates compound instructions (e.g., ”add a vase, change curtains, remove painting”).
There are also some datasets that are not commonly used, including Spot-the-Diff [228] and GeneCIS [107].
@@ 6.2.2 Results and Analysis
We evaluate the performance of compositional image retrieval methods on several benchmarks in the natural image domain. Compared to fashion retrieval, natural image tasks involve greater semantic diversity, scene complexity, and more open-ended language descriptions, posing distinct challenges for generalization and robustness.
The CSS dataset (Table IV) is one of the earliest benchmarks, constructed using a 3D rendering engine to generate synthetic scenes with precise control over object attributes, viewpoints, and modifications. This design allows for targeted evaluation of a model’s capacity to learn compositional concepts. For instance, MAAF achieves a Recall@1 of 87.8% on the 3D-to-3D subtask, demonstrating strong performance under idealized conditions. However, the simplicity of the synthetic scenes limits the dataset’s ability to reflect real-world visual complexity. Consequently, CSS is now rarely used to evaluate model performance in natural image settings.
In contrast, the CIRR dataset (Table V) has emerged as the most widely adopted benchmark for natural image retrieval with compositional queries. It emphasizes fine-grained discrimination within visually similar scenes or categories, and performance is evaluated using Recall@K (K = 1, 5, 10). Early supervised methods, such as CIRPLANT, which employed a ResNet+LSTM architecture, achieved only 15.18% Recall@1. The introduction of CLIP significantly improved retrieval performance; Combiner and CLIP4Cir achieved Recall@1 scores approaching 40%. More recent models such as SDQUR and ConText-CIR, built on stronger vision-language backbones (e.g., BLIP/BLIP2) and advanced fusion mechanisms, further elevated Recall@1 to 53–55%. Notably, SDQUR reaches 79.47% on Recall_subset@1, illustrating its capability to resolve fine-grained semantic ambiguities. Semi-supervised approaches such as VDG and CASE also show strong performance, leveraging large-scale pseudo-triplets to achieve Recall@1 of 50.96% and 48.00%, respectively, outperforming most fully supervised baselines. These results indicate that automatically constructed training data, when of sufficiently high quality, can achieve performance comparable to or exceeding that of limited human annotations in practical settings. In the zero-shot setting, models like SEIZE and PrediCIR, built on ViT-G/14, reach approximately 38% Recall@1, showing promising results without task-specific supervision, though still behind the best SL and SSL models.
The CIRCO dataset (Table VI) introduces a more challenging setup by associating each query with multiple valid target images and expanding the retrieval pool to the full COCO dataset (approximately 120K images). Evaluation is based on mAP@1, 10, 25, 50, measuring robustness to ambiguity and scalability in open-world retrieval. Results indicate a strong correlation between model capacity and performance. For example, CIReVL’s mAP@5 improves from 14.94% (ViT-B/32) to 26.77% (ViT-G/14), while LDRE increases from 23.35% (ViT-L/14) to 31.12% (ViT-G/14). This suggests that larger models generalize better to diverse target sets. Semi-supervised methods such as ConText-CIR (mAP@50 34.72%) and IP-CIR (mAP@50 38.03%) benefit from enhanced training paradigms and data augmentation strategies. Despite these gains, the absolute performance on CIRCO remains modest, with IP-CIR representing the current best at 38.03% mAP@50. This highlights ongoing challenges in fine-grained discrimination and semantic ambiguity resolution in large-scale retrieval.
In addition, Table VII presents results on two specialized datasets: MIT-States, which focuses on object-state composition, and Birds-to-Words, which targets long-form descriptive query understanding. Although these datasets are less frequently used in recent work, they remain valuable for evaluating a model’s capacity to capture nuanced semantic shifts and process complex textual inputs.
Overall, while method trends in natural image retrieval share some similarities with the fashion domain, the focus shifts more toward handling semantic complexity, open-set diversity, and one-to-many relationships. The performance on CIRR and CIRCO reveals that modeling ambiguity, leveraging high-quality supervision, and improving multimodal fusion remain critical areas for future research.
@@
CIRPLANT21’ICCV [21] ResNet152 LSTM $15.18$ $43.36$ $60.48$ $33.81$ $56.99$ $75.40$
CIRPLANT(OSCAR)21’ICCV [21] ResNet152 LSTM $19.55$ $52.55$ $68.39$ $39.20$ $63.03$ $79.49$
Combiner22’CVPR-D [33] CLIP(RN50) CLIP $33.59$ $65.35$ $77.35$ $62.39$ $81.81$ $92.02$
CLIP4Cir122’CVPR-W [34] CLIP(RN50) CLIP $38.53$ $69.98$ $81.86$ $68.19$ $85.64$ $94.17$
NEUCORE23’NIPS-W [46] ResNet BiGRU $18.46$ $49.40$ $63.57$ $44.27$ $67.06$ $78.92$
ComqueryFormer23’TMM [238] Swin-T BERT $25.76$ $61.76$ $75.90$ $51.86$ $76.26$ $89.25$
ARTEMIS22’ICLR [129] ResNet50 BiGRU $16.96$ $46.10$ $61.31$ $39.99$ $62.20$ $75.67$
RankUn23’Arxiv [100] CLIP(RN50) CLIP $32.24$ $66.63$ $79.23$ $61.25$ $81.33$ $92.02$
CLIP4Cir223’TOMM [35] CLIP(RN50) CLIP $42.05$ $76.13$ $86.51$ $70.15$ $87.18$ $94.40$
SSN23’Arxiv [67] [67] CLIP(ViT-B/32) CLIP $43.91$ $77.25$ $86.48$ $71.76$ $88.63$ $95.54$
TG-CIR23’ACMMM [72] CLIP(ViT-B/16) CLIP $45.25$ $78.29$ $87.16$ $72.84$ $89.25$ $94.13$
MCEMiner24’TIP [71] ResNet152 LSTM $17.48$ $46.13$ $62.17$
CaLa-CLIP4Cir24’SIGIR [60] $35.37$ $68.89$ $80.04$ $66.68$ $84.65$ $93.42$
CaLa-ARTEMIS24’SIGIR [60] $47.37$ $79.33$ $88.17$ $76.02$ $90.29$ $96.19$
CaLa-BLIP2cir24’SIGIR [60] $49.11$ $81.21$ $89.59$ $76.27$ $91.04$ $96.46$
BLIP4CIR24’WACV [93] BLIP BLIP $40.17$ $71.81$ $83.18$ $72.34$ $88.70$ $95.23$
BLIP4CIR+Bi24’WACV [93] BLIP BLIP $40.15$ $73.08$ $83.88$ $72.10$ $88.27$ $95.93$
SPIRIT24’TOMM [61] CLIP(RN50) CLIP $40.23$ $75.10$ $84.16$ $73.74$ $89.60$ $95.93$
DQU-CIR24’SIGIR [130] CLIP(ViT-H/14) $46.22$ $78.17$ $87.64$ $70.92$ $87.69$ $94.68$
BLIP4CIR224’TMLR [215] BLIP(ViT-B) BLIP $44.70$ $76.59$ $86.43$ $75.02$ $89.92$ $95.64$
CompoDiff(ST18M & FT)23’TMLR [98] CLIP(ViT-L/14) CLIP $21.30$ $55.01$ $72.62$ $58.82$ $77.60$ $88.37$
CompoDiff(ST18M & FT)23’TMLR [98] CLIP(ViT-G/14) CLIP $32.39$ $57.61$ $77.25$ $67.88$ $85.29$ $94.07$
SDQUR24’TCSVT [65] BLIP2 $53.13$ $83.16$ $90.60$ $79.47$ $91.74$ $96.63$
NSFSE24’TMM [47] ResNet152 BiGRU $20.70$ $52.50$ $67.96$ $44.20$ $65.53$ $78.50$
MANME24’TCSVT [53] ResNet50 BiGRU $18.27$ $48.02$ $63.23$ $42.43$ $64.89$ $77.93$
SPRC24’ICLR [63] BLIP2(ViT-L) $51.96$ $82.12$ $89.74$ $80.65$ $92.31$ $96.60$
@@
PALAVAR22’ECCV [239] CLIP(ViT-B/32) $16.62$ $43.49$ $58.51$
Pic2Word23’CVPR [12] CLIP(ViT-L/14) $23.90$ $51.70$ $65.30$
SEARLE23’ICCV [23] CLIP(ViT-B/16) $24.00$ $53.42$ $66.82$ $54.89$ $76.60$ $88.19$
SEARLE-XL-OTI23’ICCV [23] CLIP(ViT-L/14) $24.87$ $52.31$ $66.29$ $53.80$ $74.31$ $86.94$
SEARLE-XL23’ICCV [23] CLIP(ViT-L/14) $24.24$ $52.48$ $66.29$ $53.76$ $75.01$ $88.19$
MTCIR23’Arxiv [89] CLIP(ViT-L/14) $25.52$ $54.58$ $67.59$ $55.64$ $77.54$ $89.47$
KEDs24’CVPR [81] CLIP(ViT-L/14) $26.40$ $54.80$ $67.20$
PM24’Arxiv [240] CLIP(ViT-L/14) $26.10$ $55.20$ $67.50$ $56.00$ $76.60$ $88.00$
FTI4CIR24’SIGIR [79] CLIP(ViT-L/14) $25.90$ $55.61$ $67.66$ $55.21$ $75.88$ $87.98$
Denoise-I2W24’Arxiv [141] CLIP(ViT-L/14) $26.9$ $57.2$ $69.8$ $90.6$
MoTaDual(LinCIR)24’Arxiv [86] CLIP(ViT-L/14) $27.28$ $56.39$
MoTaDual(LinCIR)24’Arxiv [86] CLIP(ViT-G/14) $38.10$ $68.94$
InstructCIR124’Arxiv [241] CLIP(ViT-L/14) $35.18$ $65.12$ $77.61$ $67.54$ $84.77$
DeG25’Arxiv [140] CLIP(ViT-L/14) $26.8$ $55.0$ $67.7$
Slerp+24’Arxiv [90] BLIP $39.74$ $67.74$ $77.40$ $91.55$ $70.65$ $86.72$
TSCIR25’Arxiv [139] CLIP(ViT-L/14) $26.10$ $55.15$ $68.66$ $90.06$
SEIZE24’ACMMM [137] CLIP(ViT-L/14) $28.65$ $57.16$ $69.23$ $66.22$ $84.05$
SEIZE24’ACMMM [137] CLIP(ViT-G/14) $38.87$ $69.42$ $79.42$ $74.15$ $89.23$
PrediCIR25’CVPR [138] CLIP(ViT-L/14) $27.2$ $57.0$ $70.2$
PrediCIR25’CVPR [138] CLIP(ViT-G/14) $37.0$ $66.1$ $77.9$
CoLLM25’CVPR [142] CLIP(ViT-L/14) $29.7$ $72.8$ $91.5$
@@
MCL(OPT-6.7B)21’ICML [85] $24.15$ $55.98$ $90.92$ $59.52$
MCL(Llama2-7B)21’ICML [85] $26.22$ $56.84$ $91.35$ $61.45$
CompoDiff(ST18M)23’TMLR [98] CLIP(ViT-L/14) $18.24$ $53.14$ $70.82$ $57.42$ $77.10$ $87.90$
CompoDiff(ST18M)23’TMLR [98] CLIP(ViT-G/14) $26.71$ $55.14$ $74.52$ $64.54$ $82.39$ $91.81$
TransAgg23’BMVC [91] BLIP $38.10$ $68.42$ $93.51$
CASE24’AAAI [96] BLIP(ViT)+BERT $48.00$ $79.11$ $87.25$ $75.88$ $90.58$ $96.00$
VISTA24’ACL [73] $76.10$ $75.70$
HyCIR(CC3M+synthetic)24’Arxiv [94] CLIP $25.08$ $53.49$ $67.03$ $53.83$ $75.06$ $87.18$
HyCIR(CC3M+synthetic)24’Arxiv [94] BLIP $38.28$ $69.03$ $79.71$ $66.79$ $84.79$ $93.06$
VDG(Human+COCO’se)24’CVPR [95] BLIP $49.37$ $78.12$ $85.52$ $76.68$ $90.46$ $96.05$
VDG(Human+NLVR2’se)24’CVPR [95] BLIP $50.96$ $80.15$ $86.86$ $77.45$ $90.65$ $96.10$
PTG(SPRC+PTG)24’Arxiv [92] $36.40$ $66.10$
SCOT25’Arxiv [144] BLIP(ViT-L/16) $36.31$ $66.19$ $77.37$ $92.96$ $64.73$ $83.20$
SCOT25’Arxiv [144] BLIP2(ViT-G/14) $36.82$ $64.34$ $74.48$ $93.42$ $75.73$ $88.70$
InstructCIR225’Arxiv [143] CLIP(ViT-L/14) $50.70$ $81.61$ $98.27$ $76.10$
TSCIR25’Arxiv [139] CLIP(ViT-L/14) $29.16$ $59.33$ $71.88$ $91.52$
MRA-CIR25’Arxiv [150] BLIP2(ViT-L/14) $37.98$ $67.45$ $78.07$ $93.98$ $64.17$ $83.01$
IP-CIR(LDRE)25’CVPR [149] CLIP(ViT-L/14) $29.76$ $58.82$ $71.21$ $90.41$ $62.48$ $81.64$
IP-CIR(LDRE)25’CVPR [149] CLIP(ViT-G/14) $39.25$ $70.07$ $80.00$ $94.89$ $69.95$ $86.87$
CIG-XL(SEARLE)25’CVPR [148] CLIP(ViT-B/32) $24.75$ $54.36$ $67.81$ $90.58$ $56.24$ $77.18$
CIG-XL(LinCIR)25’CVPR [148] CLIP(ViT-L/14) $25.06$ $53.69$ $66.99$ $89.01$ $55.78$ $76.63$
ConText-CIR25’CVPR [148] CLIP(ViT-L/14) $52.65$ $83.27$ $89.51$ $98.87$ $80.32$ $92.13$
ConText-CIR25’CVPR [148] CLIP(ViT-H/14) $55.24$ $84.85$ $90.75$ $98.82$ $82.96$ $93.12$
CoLLM25’CVPR [142] CLIP(ViT-L/14) $34.7$ $77.0$ $93.1$
TABLE VI: Comparison of Different Methods on CIRCO Dataset.
Method Backbone CIRCO mAP@
Visual Textual 5 10 25 50
@@
PALAVAR22’ECCV [239] CLIP(ViT-B/32) $4.61$ $5.32$ $6.33$ $6.80$
MTCIR23’Arxiv [89] CLIP(ViT-L/14) $10.36$ $11.63$ $12.95$ $13.67$
Pic2Word23’CVPR [12] CLIP(ViT-L/14) $8.72$ $9.51$ $10.64$ $11.29$
SEARLE-XL-OTI23’ICCV [23] CLIP(ViT-L/14) $10.18$ $11.03$ $12.72$ $13.67$
SEARLE-XL23’ICCV [23] CLIP(ViT-L/14) $11.68$ $12.73$ $14.33$ $15.12$
FTI4CIR24’SIGIR [79] CLIP(ViT-L/14) $15.05$ $16.32$ $18.06$ $19.05$
CIReVL24’ICLR [74] CLIP(ViT-B/32) $14.94$ $15.42$ $17.00$ $17.82$
CIReVL24’ICLR [74] CLIP(ViT-L/14) $18.57$ $19.01$ $20.89$ $21.80$
CIReVL24’ICLR [74] CLIP(ViT-G/14) $26.77$ $27.59$ $29.96$ $31.03$
LinCIR24’CVPR [82] CLIP(ViT-L/14) $12.59$ $13.58$ $15.00$ $15.85$
LinCIR24’CVPR [82] CLIP(ViT-H/14) $17.60$ $18.52$ $20.46$ $21.39$
LinCIR24’CVPR [82] CLIP(ViT-G/14) $19.71$ $21.01$ $23.13$ $24.18$
iSEARLE-XL-OTI24’Arxiv [77] CLIP(ViT-L/14) $11.31$ $12.67$ $14.46$ $15.34$
iSEARLE-XL24’Arxiv [77] CLIP(ViT-L/14) $12.50$ $13.61$ $15.36$ $16.25$
RTD(SEARLE)24’Arxiv [83] CLIP(ViT-B/32) $11.26$ $12.11$ $13.63$ $14.37$